# Autonomous Agents for Collaborative Task under Information Asymmetry

### v1

@@ 4 Experimental Setup
We generically treat chat histories as human information. This approach simplifies modeling information asymmetry in social networks. Other types of information, such as knowledge bases, documents, or web content, can all be organized in mixed memory so iAgents is adaptable to all these kinds of information. We conduct all experiments with a maximum of 10 communication turns for agents. The experiments use GPT4 (gpt-4-0125-preview), GPT3.5 (gpt-3.5-turbo-16k), Gemini (gemini-1.0-pro-latest), and Claude (claude-sonnet) 222as of 20240501. as LLM backends. The temperature is set to 0.2. For Fuzzy Memory, we use gpt-4-0125-preview to summarize session text and OpenAI text-embedding-3-small to generate embeddings for ANN embedding search. We use precision as the metric for questions in the NP, ScheduleEasy, and FriendsTV datasets. For the ScheduleMedium and ScheduleHard datasets, we use F1 and IoU as the metrics, corresponding to the algorithm used. For the Schedule and NP datasets, we do not activate mixed memory since the information scale is small and can be fully loaded in the LLM context. Additionally, for the Schedule dataset, we do not activate the agent’s ability to initiate new communication due to the small scale of the social network.
@@ 5.1 InformativeBench Evaluations
LLM Backend Needle-Type Algorithm-Type
NP FriendsTV ScheduleEasy ScheduleMedium ScheduleHard
GPT 4 64.00% 57.94% 56.67% 51.00% 22.80%
GPT 3.5 51.00% 35.71% 36.67% 18.00% 12.25%
Claude Sonnet 50.00% 34.13% 43.33% 17.44% 18.66%
Gemini 1.0 40.00% 28.57% 26.67% 22.33% 14.40%
@@
We first comprehensively assessed the performance of iAgents using some state-of-the-art LLMs on InformativeBench, as shown in Table 3. GPT-4 achieves over 50% accuracy across most datasets, indicating its potential to work on behalf of humans for cooperation. However, smaller-scale LLMs still face significant challenges in solving cooperation problems in information asymmetry. The NP dataset is the simplest, requiring agents to find relevant information within the context and exchange opinions. Most models could only achieve about 50% precision on the easiest NP task. This is consistent with recent research findings hsieh2024ruler that simply locating information is insufficient to assess a model’s ability to analyze in-context/long-context information. In the NP dataset, agents are required to locate information within a relatively short context length, but the information is dispersed between two parties, creating an asymmetry. This remains a challenging task for current LLM agents. For the Schedule dataset, agents must locate information and enhance their communication skills for further reasoning and computation. As questions become harder, performance drops, with most models solving less than 20% of the hardest questions. The FriendsTV dataset introduces a large social network, requiring agents to use external memory to retrieve rationale from extensive human information. Most LLMs struggle to exceed 40% accuracy in this dataset. Thus, while previous studies show impressive performance when agents are omniscient, collaborating in information asymmetry remains challenging.
@@ 5.2 Ablation Study
Figure 7: Ablation study on a) InfoNav and b) other mechanisms including Distinct Memory, Fuzzy Memory, and Recursive Communication. Experiments are conducted using GPT 3.5 as the backend.
We conducted ablation experiments on several key designs of the iAgents framework, as detailed in Figure 7. Analyzing the FriendsTV dataset revealed that incorporation of the mixed memory mechanism led to a performance increase ranging from 2.38% to 6.34%, surpassing the impact of InfoNav, which resulted in only a 0.8% performance increase. This discrepancy underscores the greater significance of effective retrieval over reasoning during communication in large social networks with mass information. Notably, the ablation of both memory mechanisms emphasized the indispensability of mixed memory. The introduction of recursive communication exhibited the most significant performance gain (12.7%), primarily due to the challenges posed by the vast social network in the FriendsTV dataset. By actively introducing new communications within ongoing dialogues, agents could acquire and corroborate information, thus significantly enhancing performance. This highlights the imperative of scalability in our proposed framework for addressing real-world problems.
@@ 5.3 Analysis on Agents’ Behaviour
Figure 8: The figure depicts the distribution of different behaviors of agents in adjusting memory retrieval based on the progress of communication. Different colors denote different behaviors such as maintaining, increasing, or decreasing parameters, while different textures indicate whether the agent ultimately completes the task. Agents predominantly tend to maintain parameters unchanged, but when changes occur, they tend to increase parameters to gain more information.
Memory Behaviour Similarly, we explored how agents adapt their memory retrieval strategies during communication. We examined three parameters in distinct memory queries: the context window, which determines the breadth of contextual messages; the total message retrieval limit; and the size of the query keywords set. For fuzzy memory, we analyzed two parameters: the number of queried responses (topk) and the length of the query text. These findings are illustrated in Figure 8. Our analysis revealed several notable trends. The majority of agents do not change their behavior during communication. However, when agents decide to change their behavior, we observed that they tended to increase the amount of retrieved information over time. This augmentation trend was particularly pronounced on the overall message retrieval limit, where the frequency of increase actions surpassed that of decrease actions by nearly threefold. Furthermore, agents who completed tasks exhibited a more conservative approach, with a lower proportion of behavioral changes compared to agents unable to complete tasks. This phenomenon may be attributed to the difficulty of certain tasks, making agents continuously refine their strategies in pursuit of the required information.
@@ 5.4 Analysis on Real World Concern
Study Experiments FriendsTV
Base iAgents 35.71%
Prior Knowledge Distraction Anonymous 32.54%
Privacy Study PrivacyPrompt 30.95%
Table 5: Analysis experiments on prior knowledge and privacy. Use GPT3.5 as LLM backend.
We studied two significant challenges in extending the iAgents to real-world applications, as shown in Table 5. Firstly, we investigated whether the agent can effectively respond to human input without being overly influenced by factual knowledge obtained during pre-training wu2024faithful ; qi2024can . Secondly, we explored the agent’s ability to engage in communication while upholding human privacy. Our experiments were conducted using the GPT3.5 model on the FriendsTV dataset.
@@
Privacy Concern In investigating whether agents can communicate without compromising privacy, we conducted an experiment involving modifications to the agent’s system prompt, emphasizing the importance of privacy preservation in utterances. The agent then utilized vague expressions such as "somebody/somewhere" and disclosed only entity information relevant to the task. This adjustment led to a performance drop from 35.71% to 30.95%, indicating the ongoing challenge of achieving collaboration while ensuring privacy. It’s important to note that we solely adjusted privacy settings on the output side, rather than restricting agent access to human information on the input side. This decision was made because setting access permissions might inadvertently reveal prior task-related information. Thus, the real challenge lies in appropriately regulating access to information based on task requirements, akin to teaching the agent to retrieve necessary information accurately. What’s more, human user can customize their personal file permissions in real-life applications, which is more like an engineering issue. Lastly, absolute privacy protection is impractical, as absolute privacy protection amounts to forgoing problem-solving through collaboration.
6 Related Work
LLM Agents Originating from ancient Greek philosophy, the concept of an "agent" referred to a being that possesses the capacity to act with intentionality, often driven by the manifestation of certain mental states and events, such as desires, beliefs, and intentions sep-agency . As AI evolves, the "agent" concept is incorporated to facilitate the simulation and comprehension of intelligent behavior. In this context, an agent is typically defined as a computer system that is autonomous, interactive, reactive, and proactive Wooldridge_Jennings_1995 . Prior to the advent of Large Language Models (LLM), most agent research concentrated on augmenting particular abilities, such as symbolic reasoning, or excelling in specific tasks like Go 10.1109/64.180407  10.1145/176789.176804  sutton1999reinforcement . However, with the emergence of LLM, the focus of agent research has shifted. In the evaluation by  bubeck2023sparks , GPT-4 is considered to have achieved a form of general intelligence, thus equipping LLMs with agency and intrinsic motivation is an intriguing and important direction. Building on this,  xi2023rise introduces a framework for LLM-based agents, encompassing a brain for decision-making, a perception module for sensory input, and an action module for environmental interaction. A growing number of studies have begun to utilize LLM as a primary component of the brain to construct artificial intelligence agents weng2023agent , and apply them in various real-world scenarios. This is largely due to the inherent characteristics of LLMs, such as their demonstrated autonomy, reactivity, pro-activeness, and social ability, which align well with the key attributes of an agent xi2023rise . For instance,  qian2023communicative creates a virtual software development company using LLM-based agents, who collaborate through a chat chain to break down, propose, and validate solutions for atomic subtasks, thereby efficiently completing the entire software development process.  boiko2023emergent presents an agent system designed for conducting scientific experiments. However, these studies primarily focus on the capabilities of agents, overlooking the interaction and cooperation paradigms of agents as conscious entities. This research gap warrants further exploration.
Paradigms of Human-Agent and Multi-Agent Cooperation To ensure the actions of agents align with human objectives kenton2021alignment , uphold the safety, legality, and morality of agent behaviors, as well as address data privacy issues and compensate for data scarcity in specific domains paul2023 , human-agent interaction is indispensable. Currently, there are two main paradigms of human-agent interaction. The first, the Equal Partnership Paradigm, views agents as communicators who understand human emotions and collaborate from a human perspective Hasan_2023 . The second, the Instructor-Executor Paradigm, emphasizes the human’s guiding role, with agents interpreting and executing human instructions gao2023assistgpt . Although LLM-based agents possess exceptional capabilities in solving complex tasks, they exhibit significant limitations when operating as isolated entities. Specifically, single LLM-based agents are limited by their inability to collaborate, learn from social interactions and multi-turn feedback liu2023training , and function effectively in complex, multi-agent scenarios xi2023rise . Many studies have shown that the interaction of multiple single agents minsky1988society , each with specific functions, can stimulate stronger intelligence Balaji2010AnIT .On one hand, allowing multiple intelligent agents to collaborate can handle complex tasks more efficiently. For instance, based on multi-agent cooperation,  mandi2023roco utilizes multi-agent systems to collectively reason about task strategies, and decompose tasks during planning to accelerate trajectory planning. On the other hand, introducing game theory concepts into multi-agent systems article , where each agent adjusts its strategy based on the behaviors and potential responses of other agents in a competitive environment, can lead to more powerful behaviors liang2023encouraging . Some of the latest research has begun to focus on multi-agent simulation scenarios where there exists information asymmetry zhou2024sotopia ; zhou2024real . They found that the dialogue participants simulated in an omniscient environment were more successful in achieving social goals than those in a non-omniscient environment, despite the latter being closer to reality. In addition, they also demonstrated that learning from omniscient simulations can enhance the naturalness of interactions, but it hardly improves the achievement of goals in cooperative scenarios. Most real-world application scenarios involve information asymmetry. Therefore, how to enable agents, each possessing private user information, to cooperate in assisting humans in problem-solving, is a necessary issue to be addressed for the real-world application of agents in human society.
LLM Agent Reasoning In previous research, agents and LLMs have been tasked with providing accurate information to human users in human-machine collaboration. Due to the nature of language models, the output of information relies on the user’s input and the previously decoded content serving as rationale context. Therefore, some work on reasoning has explored how to improve the organization and expression of this rationale context NEURIPS2022_9d560961 ; NEURIPS2023_271db992 ; besta2024graph ; ding2023everything , to enhance the accuracy of output information. However, some research has also found that the reasoning process of LLMs differs from that of humans yang2024large ; jin2024impact ; chen2024masked ; pfau2024let . For questions relying on internal knowledge within LLMs to answer, agents do not necessarily solve problems step by step like humans. Instead, compared to steps, having context with sufficient information content is more important. In this paper, we focus on machine-to-machine communication, relying on external information of LLMs to collaboratively answer questions. This poses different requirements for LLM reasoning, especially in terms of how to promote information flow so sufficient information is included in the context provided to LLM. Agents need to actively hu2024uncertainty and accurately acquire, provide, and ask for information.

### v2

@@ 5 Experimental Setup
We generically treat chat histories as human information. This approach simplifies modeling information asymmetry in social networks. Other types of information, such as knowledge bases, documents, or web content, can all be organized in mixed memory so iAgents is adaptable to all these kinds of information. We conduct all experiments with a maximum of 10 communication turns for agents. The experiments use gpt-4-0125-preview, gpt-3.5-turbo-16k, gemini-1.0-pro-latest, and claude-sonnet 222as of 20240501. as LLM backends. The temperature is set to 0.2. For Fuzzy Memory, we use gpt-4-0125-preview to summarize session text and OpenAI text-embedding-3-small to generate embeddings for ANN embedding search. We use precision as the metric for questions in the NP, ScheduleEasy, and FriendsTV datasets. For the ScheduleMedium and ScheduleHard datasets, we use F1 and IoU as the metrics, corresponding to the algorithm used. Details about the metrics are shown in Section C.3. For the Schedule and NP datasets, we do not activate mixed memory since the information scale is small and can be fully loaded in the LLM context. Additionally, for the Schedule dataset, we do not activate the agent’s ability to initiate new communication due to the small scale of the social network.
@@ 6.1 InformativeBench Evaluations
LLM Backend Reasoning-Oriented (Schedule Dataset) Needle-Oriented
Easy Medium Hard NP FriendsTV
GPT 4 56.67% 51.00% 22.80% 64.00% 57.94%
GPT 3.5 36.67% 18.00% 12.25% 51.00% 35.71%
Claude Sonnet 43.33% 17.44% 18.66% 50.00% 34.13%
Gemini 1.0 26.67% 22.33% 14.40% 40.00% 28.57%
@@
We first comprehensively assessed the performance of iAgents using some state-of-the-art LLMs on InformativeBench, as shown in Table 1. GPT-4 achieves over 50% accuracy across most datasets, indicating its potential to work on behalf of humans for cooperation. However, smaller-scale LLMs still face significant challenges in solving cooperation problems in information asymmetry. Most models could only achieve about 50% precision on the easiest NP task. For the Schedule dataset, as questions become harder, performance drops, with most models solving less than 20% of the hardest questions. The FriendsTV dataset introduces a large social network, requiring agents to use external memory to retrieve rationale from extensive human information. Most LLMs struggle to exceed 40% accuracy in this dataset. Thus, while previous studies show impressive performance when agents are omniscient, collaborating in information asymmetry remains challenging.
@@ 6.2 Ablation Study
Experiment Reasoning-Oriented (Schedule Dataset) Needle-Oriented
Easy Medium Hard NP FriendsTV
iAgents (Full Model) 36.67% 18.00% 12.25% 51.00% 35.71%
Ablation on InfoNav:
w/o InfoNav 10.00% 3.56% 7.34% 39.00% 34.92%
Ablation on other mechanisms (Limited Applicability):
w/o Recursive Comm 48.00% 23.02%
w/o Fuzzy Memory 29.37%
w/o Clear Memory 33.33%
Table 2: Ablation study on iAgents. Dashes (–) indicate: (1) iAgents on Reasoning-Oriented dataset does not equip other mechanisms, hence no ablation needed; (2) For NP dataset, iAgents does not utilize Mixed Memory hence there is no ablation.
We conducted ablation experiments on several key designs of the iAgents framework, as detailed in Table 2. Analyzing the FriendsTV dataset revealed that incorporation of the mixed memory mechanism led to a performance increase ranging from 2.38% to 6.34%, surpassing the impact of InfoNav, which resulted in only a 0.8% performance increase. This discrepancy underscores the greater significance of effective retrieval over reasoning during communication in large social networks with mass information. Notably, the ablation of both memory mechanisms emphasized the indispensability of mixed memory. The introduction of recursive communication exhibited the most significant performance gain (12.7%), primarily due to the challenges posed by the vast social network in the FriendsTV dataset. By actively introducing new communications within ongoing dialogues, agents could acquire and corroborate information, thus significantly enhancing performance. This highlights the imperative of scalability in our proposed framework for addressing real-world problems.
@@ 6.3 Analysis on Agents’ Behaviour
Figure 5: The figure depicts the distribution of different behaviors of agents in adjusting memory retrieval based on the progress of communication. Agents predominantly tend to maintain parameters unchanged, but when changes occur, they tend to increase parameters to gain more information.
Memory Behaviour Similarly, we explored how agents adapt their memory retrieval strategies during communication. We examined three parameters in clear memory queries: the context window, which determines the breadth of contextual messages; the total message retrieval limit; and the size of the query keywords set. For fuzzy memory, we analyzed two parameters: the number of queried responses (topk) and the length of the query text. These findings are illustrated in Figure 5. Our analysis revealed several notable trends. The majority of agents do not change their behavior during communication. However, when agents decide to change their behavior, we observed that they tended to increase the amount of retrieved information over time. This augmentation trend was particularly pronounced on the overall message retrieval limit, where the frequency of “increase” actions surpassed that of “decrease” actions by nearly threefold. Furthermore, agents who completed tasks exhibited a more conservative approach, with a lower proportion of behavioral changes compared to agents unable to complete tasks. This phenomenon may be attributed to the difficulty of certain tasks, making agents continuously refine their strategies in pursuit of the required information.
@@ 6.4 Analysis on Real World Concern
We studied two significant challenges in extending the iAgents to real-world applications. Firstly, we investigated whether the agent can effectively respond to human input without being overly influenced by factual knowledge obtained during pre-training wu2024faithful ; qi2024can . Secondly, we explored the agent’s ability to engage in communication while upholding human privacy. Our experiments were conducted using the GPT3.5 model on the FriendsTV dataset.
@@
Privacy Concern In investigating whether agents can communicate without compromising privacy, we conducted an experiment involving modifications to the agent’s system prompt, emphasizing the importance of privacy preservation in utterances. The agent then utilized vague expressions such as "somebody/somewhere" and disclosed only relevant entity information. This adjustment led to a performance drop from 35.71% to 30.95%, indicating the ongoing challenge of achieving collaboration while ensuring privacy. It’s important to note that we solely adjusted privacy settings on the output side, rather than restricting agent access to human information on the input side. This decision was made because setting access permissions might inadvertently reveal prior task-related information. Thus, the real challenge lies in appropriately regulating access to information based on task requirements, akin to teaching the agent to retrieve necessary information accurately. Additionally, absolute privacy protection is impractical, as absolute privacy protection amounts to forgoing problem-solving through collaboration.
7 Limitations
While the iAgents framework introduces innovative multi-agent collaboration, it has several limitations and challenges. Privacy Issues: As discussed in Section 6.4, we examined the performance of agents communicating and collaborating under privacy constraints, highlighting the trade-off between privacy and collaboration. We define three privacy levels. L1: Users fully share personal information, allowing maximum efficiency for iAgents. L2: Users keep their personal information private. iAgents can handle this situation by deploying an edge-side small language model agent for information acquisition. L3: Users demand maximal privacy, with both personal and agent communication handled locally on private devices, which is still a challenge for small language model agents. Network Modeling: The current framework initiates communications based on user relevance but lacks nuanced modeling of human social networks and collaboration history. Enhancing network topology through added or removed nodes and incorporating past interactions could improve communication efficiency. Human-Agent Interaction: Although iAgents target full autonomy, human involvement for verification remains necessary in real-world scenarios. Processes must be designed to prompt user feedback and adjust agent strategies accordingly, ensuring alignment with user preferences. Cost: The high input token consumption (about 30,000 tokens per task) required for iAgents to handle human information is a challenge. However, advancements in long-context models, which extend input token length, present opportunities to reduce the cost of scaling iAgents.